OVERTONE AND COMBINATION BAND SPECTROSCOPY OF JET COOLED METHANOL

DAVID RUEDA, O.V. BOYARKIN, AND T.R. RIZZO, INSTITUT DE CHIMIE PHYSIQUE MOLECULAIRE, EPFL, CH - 1015 LAUSANNE; ANDREI CHIROKOLAVA, INDRANATH MUKHOPADHYAY, AND \underline{DAVID S PERRY}, DEPARTMENT OF CHEMISTRY, UNIVERSITY OF AKRON, AKRON, OHIO 44325.

Overtone spectra of jet-cooled methanol have been recorded from 5,000 to 14,000 cm^{−1} using Infrared Laser Assisted Photofragment Spectroscopy (IRLAPS) for the detection of the vibrationally excited molecules. In addition to the OH stretch overtones (nν_{1}), the major components of the spectra are the overtones of the CH stretch (up to 5ν_{CH}) as well as combinations of the OH stretch with the CO stretch (nν_{1}+ν_{8}), the COH bend (nν_{1}+ν_{6}), and both (nν_{1}+ν_{6}+ν_{8}). These data, together with photoacoustic data [1] up to 18,250 cm^{−1} and previously reported IRLAPS data [2] up to 22,166 cm^{−1}, have been fit to an anharmonic Hamiltonian. In this Hamiltonian, the CH stretch vibrations are treated as a pair of local modes, ν_{a} and ν_{b}, for v_{CH} = 3 and higher. The other modes are treated as the usual normal vibrations.

[1] H.L. Fang, D.M. Meister, R.L. Swofford, J. Phys. Chem. 88, 405-409 (1984).

[2] O.V. Boyarkin, L. Lubich, R.D.F. Settle, D.S. Perry, T.R. Rizzo, J. Chem. Phys. 107, 8409-8422 (1997).